If f(x)=the square root of (x^2-9), then f(4) is?
Answer by stanbon(75887)

If f(x)=the square root of (x^2-9), then f(4) is?
----
f(x) = sqrt(x^2-9)
----
Ans: f(4) = sqrt(16-9) = sqrt(7)
----------------
Cheers,
Stan H.
----------------